OpenStudy (anonymous):

The coefficient of static friction for between a 9-newton wooden blok an a level surface is 0.4 and the coefficient of kinetic friction is 0.3. Determine the force of friction, on the block if you pull it horizontally with the follow force. Specify if the friction is static or kinetic: (a) 0-N (b)2-N (c)4-N (d)6-N (e)8-N how do i do this

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

Fg=9N µs=0.4 µk=0.3 Ff=?  Ff is given by the formula \(\Large Ff=\frac{\mu}{F_n}\)

OpenStudy (anonymous):

oops made a mistake! it should be \(F_f = \mu \times F_n\)

OpenStudy (anonymous):

so i have 0.3*9=2.7

OpenStudy (anonymous):

what does it mean when the question says "with the follow force" (i just want to make sure that i'm telling you the right method)

OpenStudy (anonymous):

i'm not sure maybe with the following force they mean 9-newton

OpenStudy (anonymous):

oh i should get 29.4 not 2.7

OpenStudy (anonymous):

okay. Since the object is \(resting\) on a level surface before you pull it, you should use the \(maximum\) coefficient of static friction and not the coefficient of kinetic friction because you have to overcome the \(maximum\) coefficient of static friction before making the block move

OpenStudy (anonymous):

so instead of using 0.3 as your \(\mu\), you should use 0.4 because it is the \(\mu _{max}\)

OpenStudy (anonymous):

*\(\Large \mu _{s_{max}}\)

OpenStudy (anonymous):

okay so when i do that i should have .4*98=39.2

OpenStudy (anonymous):

where did you get 98? Fg=9N Fg=Fn therefore Fn=9N

OpenStudy (anonymous):

gravity?

OpenStudy (anonymous):

soi should do .4*9=3.6

OpenStudy (anonymous):

Fn is the normal force Fg is the force of gravity gravitational field strength(g) is different from the force of gravity(Fg) Fg=mg but since the question already provided you the weight of the block or the Fg, you dont have to calculate for Fg and no need to use the value of g. Yes that will be right. Round your answer and you'll find it in one of your choices :) Gtg , goodnight!

OpenStudy (anonymous):

thanks